OpenStudy (anonymous):

Divide and simplify. 7/3 / (divide) 9/13

OpenStudy (ivancsc1996):

\[\frac{ \frac{ 7 }{ 3 } }{ \frac{ 9 }{ 13 } }=\frac{ 7 \times 13 }{ 9 \times 3 }= \frac{ 91 }{ 27 }\] I don't think you can simplify that anymore.

OpenStudy (anonymous):

A better way is to know that dividing a fraction is exactly like multiplying by its recipical.